# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ 2 Canonical Constant-Roll Inflation
Figure 2 shows all constant-roll potentials (2.17), (2.21), (2.25), and (2.32). As can be seen, the potentials are extremely flat for $\beta\rightarrow-3$ (USR limit) and $\beta\rightarrow 0$ (slow-roll limit). But when $\beta$ is exactly zero, the potential has a parabolic shape. It is true that in the case of the slow-roll, as in the case of USR, the potential should be very flat to ensure a slow rate of roll and a minimum of 40-60 e-folds of inflation. But if one defines USR as a model with a flat potential, then $\beta\rightarrow 0$ is also an USR case, and there is no clear distinction between USR and slow-roll models. Therefore, in order to avoid ambiguities, we define USR limit with condition $V_{,\phi}\ll\{\dot{\phi},\ddot{\phi}\}$, and slow-roll limit with $\{\epsilon_{1},\epsilon_{2},\text{...}\}\ll 1$ (which in the canonical case leads to $\ddot{\phi}\ll\dot{\phi}\ll 1$). This definition still allows for a model that is both USR and slow-roll (for $V_{,\phi}\ll\ddot{\phi}\ll\dot{\phi}\ll 1$). But by expanding the potential and field for solutions (2.13), (2.14), and (2.15), around 0 and $-3$, we find that for $\beta\rightarrow 0$ we have $V_{,\phi}/\ddot{\phi}\propto 1/\beta$, while for $\beta\rightarrow-3$ we have $V_{,\phi}/\ddot{\phi}\propto\beta+3$ and $V_{,\phi}/\dot{\phi}\propto\beta+3$. Thus, $\beta\rightarrow 0$ does not represent an USR case but $\beta\rightarrow-3$ does. Solution (2.29) also is not an USR case as it leads to $\ddot{\phi}=0$ and condition $V_{,\phi}\ll\ddot{\phi}$ cannot be satisfied.
It could also be argued that the name ”USR” is inappropriate. Because USR begins with a period of fast roll inflation. Nevertheless, we do not obsess over nomenclature in this work and respect the names that have long been used.
Additionally, it should be noted that all USR cases cannot be explained by constant-roll condition (2.6). In other words, any potential that is extremely flat might be considered an USR potential, and not only the constant-roll potentials described in equation (2.17). Therefore, by setting $\beta\approx-3$, we are examining only a particular set of potentials from a broader class of USR potentials. A similar argument can be made for the slow-roll limit $\beta\approx 0$. In the slow-roll approximation $\ddot{\phi}/H\dot{\phi}$ is very small and not necessarily constant.
@@
The 2018 Planck data [74] constrain the spectral index as $n_{s}=0.9649\pm 0.0042$. Accordingly, the only feasible values for the constant-roll parameter are $\beta\approx 0.018$ and $\beta\approx-3.018$. It is imperative to note that the power (2.35) is evaluated at horizon exit. However, as we will show later, in the case $\beta<-3/2$ the modes continue to grow on the supperhorizon scales and the effect of growing modes should also be taken into account in the evaluation of the power spectrum. Consequently, for $\beta\approx-3.018$, result (2.35) is invalid, and $\beta\approx 0.018$ is the only compatible point.
@@
To agree with the Planck data, $\epsilon_{1}$ should be in the range of $0.0077$ to $0.0098$, which is between $N=-65$ and $-51$. That means this case is transient and can last up to 14 e-folds. As far as we know, results (2.36) and (2.37) are new.
@@
For solution (2.31), $\beta$ is taken to be exactly zero. In this case, a slight change in initial values (that changes $\beta$) will result in switching to another constant-roll solution and never returning to $\beta=0$ again. As a result, this solution is also unstable. In general, all $\beta<-3/2$ solutions are unstable, while all $\beta\geq-3/2$ solutions, except for $\beta=0$, are stable.
This work does not focus on studying non-Gaussianity. However, the non-Gaussian properties of solution (2.13) are particularly interesting and worth discussion. In this case the contribution due to the field redefinition to the non-Gaussianity can be derived as follows (see Appendix A):
@@
A look at the evolution of curvature perturbations on supper Hubble scales ($k\ll aH$) can help explain the sudden change of behavior in equations (2.35) and (2.38) at $\beta=-3/2$. The $k^{2}$ term in the Mukhanov-Sasaki equation (A.2) is negligible at supper Hubble scales. As a result, we can integrate this equation to get [10]:
$u=C_{1}+C_{2}\int\frac{\mathrm{~{}d}\tau}{a^{2}\epsilon_{1}}=C_{1}+\frac{C_{2}}{a_{0}^{3}H}\int\frac{\mathrm{~{}d}N}{e^{3N}\epsilon_{1}},$ (2.39)
where $C_{1}$ and $C_{1}$ are two arbitrary constants, and for the last equality we have used the the relation $a=a_{0}\text{exp}(N)$ (derived by integrating the definition $\mathrm{~{}d}N\equiv\mathrm{d}\ln a$), as well as the fact that Hubble parameter is roughly constant at large scales. As can be seen in equation (2.39), the curvature perturbation has a constant part and an integral part. If $\epsilon_{1}$ is decaying slower than $e^{-3N}$, then the mode function will converge to a constant at super Hubble scales. In contrast, when $\epsilon_{1}$ decays faster than $e^{-3N}$, the integral part diverges and the mode function continues to evolve at large scales. Consistency relation requires that the long modes are well frozen out at large scales, which is not the case here. Therefore, Maldacena’s methodology for proving non-Gaussianity is small does not apply here, and consistency relation cannot be satisfied for fast decaying $\epsilon_{1}$. In equation (2.19), it is evident that for $\beta=-3/2$, the slow-roll parameter decays as $e^{-3N}$. This completely explains why equations (2.35) and (2.38) suddenly change their behavior at $\beta=-3/2$.
@@ 3 Constant-Roll Condition in Non-Canonical Models
In previous section, the constant-roll condition was defined by three equivalent expressions given in (2.6). In this section, we consider each expression as a distinct constant-roll condition and impose it on a general $P(X)$ model to show that the nature of the model and its consequences is sensitive to the expression we use. Furthermore, the equivalence is restricted to some models, like single field canonical case, and the result would be different for either expression, in other inflationary models.
A general $P(X)$ model is described by the following action,
$S=\frac{1}{2}\int d^{4}x\sqrt{-g}[R+2P(X,\phi)].$ (3.1)
In FLRW background with (,+,+,+)(-,+,+,+)( - , + , + , + ) signature, $X=\dot{\phi}^{2}/2$ corresponds to the canonical kinetic term. The Friedman equations for this action are
@@
Here $c_{s}$ represents the sound speed which can be expressed as
@@
and the non-trivial field dynamics is controlled by the friction term and an effective potential given by
@@
where subscript F refers to a constant-roll condition that constraints the field rate of roll. We call this constraint the field-constant-roll (FCR) condition. This constant-roll condition can be written as $\mathrm{~{}d}\ln\dot{\phi}/\mathrm{~{}d}N=\beta_{\text{\tiny{F}}}$, indicating that the scalar field has a constant roll rate. Integrating it out, we find
@@
As the second approach, we define a constant-roll condition that (similar to the canonical case) leads to a constant $\epsilon_{2}$. Provided that the first slow-roll parameter is small [^2] This is usual in the standard models of inflation based on an inflationary attractor dynamics., then the second slow-roll parameter $\epsilon_{2}=2\epsilon_{1}+\ddot{H}/(H\dot{H})$ is roughly constant by imposing the following condition:
@@
We refer to this constraint as the Hubble-constant-roll (HCR) condition. Note that, unlike the canonical case, $\beta_{\text{\tiny{H}}}$ here does not always equal to $\ddot{\phi}/(H\dot{\phi})$. The HCR condition does not ensure a constant rate of roll for the scalar field. Likewise, the FCR condition does not guarantee a constant $\epsilon_{2}$.
An interesting property of the HCR condition is that we can directly integrate it to obtain the Hubble parameters given by (2.13), (2.14), (2.15), and (2.29). Accordingly, the slow-roll parameters and the acceptable ranges for $N$ and $\beta_{\text{\tiny{H}}}$ are also similar to the canonical case.
@@
where we have defined the effective kinetic term, $Q(X,\phi)\equiv P(X,\phi)+V(\phi)$. From this we notice that when the FCR condition is imposed, $\ddot{H}/(H\dot{H})$ is not necessarily constant. In a particular model, if the right hand side of equation (3.10) becomes constant, the HCR condition is met. In this case we can use (3.9) to rewrite (3.10) as
@@
A key property of the canonical constant-roll model is that it leads to USR for $\beta\rightarrow-3$, and it can be considered a generalization of USR model. However, our FCR and HCR conditions do not always correspond to the generalization of USR. In other words, in a particular model, these constant-roll conditions may not reduce to a flat USR potential for any choice of the constant-roll parameter. As our final approach to the generalized constant-roll condition, we define the constant-roll conditions so that the profile of the scalar potential $V(\phi)$ is sufficiently flat when constant-roll parameter approaches $-3$. We approximate the field equation (3.4) in the USR limit as
@@
Although our third manipulation of constant-roll constraint makes implicit use of the third expression in (2.6), it is important to emphasize that condition (3.14) is not the only possible way to generalize the USR. For instance, the condition $\dot{V}=(3+\beta_{\text{\tiny{V}}})HV$, which again exhibits USR when $\beta_{\text{\tiny{V}}}\rightarrow-3$ can be viewed as another generalization of USR and may provide new insights into the model. Nevertheless, the form (3.14) has two advantages: it is identical to FCR and HCR conditions in the canonical limit, and for the examples we will discuss later, $\beta_{\text{\tiny{V}}}\rightarrow 0$ leads to slow-roll case. In this work, we discuss only a small number of an apparently endless variety of possible constant-roll conditions picked out by the demand of having canonical counterpart.
@@
To find the relation between $\beta_{\text{\tiny{V}}}$ and $\beta_{\text{\tiny{H}}}$, we elaborate on equation (3.3) to rewrite the PCR condition as
@@
This implies that given a non-canonical action, two constant-roll models based on PCR and HCR constraint would be homologous if we have
@@
Finally, note that the system of equations (3.11), (3.15), and (3.17) are dependent, in the sense that when every pair of the equations are satisfied, all three kinds of constant-roll conditions lead to similar results. The canonical constant-roll model satisfies these conditions for $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}$. In what comes next, we arrange to clearly distinguish between the above approaches for generalizing the constant-roll conditions.
@@ 4 Constant-Roll and Propagation Inflation
This section considers a special class of inflationary models with constant sound speed to demonstrate the applications of the constant-roll conditions discussed in section 3. With constant $c_{s}$, integration of equation (3.5) gives us
@@
Using relation (4.1) the dynamics of these models is governed by
@@
$\ddot{\phi}+3c_{s}^{2}H\dot{\phi}+\mathcalboondox{V}_{,\phi}=0,$ (4.4)
where derivative of effective potential is given by
@@
In this section we study three classes of constant-roll and propagation models constrained with constant-roll conditions introduced in previous section. However, before proceeding, it is useful to determine whether these classes would be homologous or not. The effective kinetic term in constant propagation model is given by
@@
Using this, we can show that constant propagation models satisfy equation (3.17) with $2\beta_{\text{\tiny{H}}}=\left(1+c_{s}^{2}\right)\beta_{\text{\tiny{V}}}$. In other words, in a constant propagation model if $\beta_{\text{\tiny{H}}}$ is constant, $\beta_{\text{\tiny{V}}}$ is also constant and vice versa. Therefore, HCR and PCR constrained classes are not independent and do not lead to different solutions. By knowing that, we can easily discuss the model’s USR limit. Since for HCR condition the Hubble parameter can be derived identical to the canonical case, and since in the constant propagation model the HCR and PCR condition are homologous, the Hubble parameter and slow-roll parameters in USR case of the constant propagation models are given by equations (2.13) and (2.19), but with $\beta$ replaced with $-3\left(c_{s}^{2}+1\right)/2$. As we discussed in section 2, if $\epsilon_{1}$ decays faster than $e^{-3N}$ modes keep growing at superhorizon scales. Thus, the model cannot be stable unless the condition $\beta=-3\left(c_{s}^{2}+1\right)/2>-3/2$ or equivalently $c_{s}^{2}<0$ is satisfied, which is impossible. It follows that the USR limit of constant propagation model (4.1) can never be stable.
@@
This means that FCR and HCR constrained classes are generally independent and lead to different solutions unless the offending non-universal term ˙$\dot{\mathcal{F}}/(2H\mathcal{F})$ is a constant.

### v3

@@ 2 Canonical Constant-Roll Inflation
Figure 2 shows all constant-roll potentials (2.17), (2.21), (2.25), and (2.32). As can be seen, the potentials are extremely flat for $\beta\rightarrow-3$ (USR limit) and $\beta\rightarrow 0$ (slow-roll limit). But when $\beta$ is exactly zero, the potential has a parabolic shape. It is true that in the case of the slow-roll, as in the case of USR, the potential should be very flat to ensure a slow rate of roll and a minimum of 40-60 $e$-folds of inflation. But if one defines USR as a model with a flat potential, then $\beta\rightarrow 0$ will be an USR case as well, and there is no clear distinction between USR and slow-roll models. Therefore, in order to avoid ambiguities, we define USR limit with condition $V_{,\phi}\ll\{H\dot{\phi},\ddot{\phi}\}$, and slow-roll limit with $\{\epsilon_{1},\epsilon_{2},\text{...}\}\ll 1$ (which in the canonical case leads to $\ddot{\phi}\ll H\dot{\phi}\ll H^{2}$). By expanding the potential and field for solutions (2.13), (2.14), and (2.15) around 0 and $-3$, we find that when $\beta\rightarrow 0$ we have $\ddot{\phi}/V_{,\phi}\propto\beta$ $V_{,\phi}/\ddot{\phi}\propto 1/\beta$, while when $\beta\rightarrow-3$ we have $V_{,\phi}/\ddot{\phi}\propto\beta+3$ and $V_{,\phi}/\dot{\phi}\propto\beta+3$. Thus, $\beta\rightarrow 0$ does not represent an USR case but $\beta\rightarrow-3$ does. Solution (2.29) also is not an USR case as it leads to $\ddot{\phi}=0$ and condition $V_{,\phi}\ll\ddot{\phi}$ cannot be satisfied.
It could also be argued that the name ”ultra slow-roll” is inappropriate. Because USR begins with a period of fast roll inflation. Nevertheless, we do not obsess over nomenclature in this work and respect the names that have long been used. Additionally, it should be noted that the model described by equations (2.16)-(2.18) is an USR model for parameter value $\beta=-3$, but the reverse is not always true. A similar argument can be made for the slow-roll limit $\beta\approx 0$. In the slow-roll approximation $\ddot{\phi}/H\dot{\phi}$ is very small and not necessarily constant.
@@
The 2018 Planck data [76] constrain the spectral index as $n_{s}=0.9649\pm 0.0042$. Accordingly, the only feasible values for the constant-roll parameter are $\beta\approx 0.018$ and $\beta\approx-3.018$. It is imperative to note that for deriving (2.35) the power is evaluated at horizon exit. However, as we will show later, in the case $\beta<-3/2$ the modes continue to grow on the supperhorizon scales and the effect of growing modes should also be taken into account in the evaluation of the power spectrum. Consequently, for $\beta\approx-3.018$, result (2.35) is invalid, and $\beta\approx 0.018$ is the only compatible point.
@@
To agree with the Planck data, $\epsilon_{1}$ should be in the range of $0.0077$ to $0.0098$, which is between $N=-65$ and $-51$. That means this solution cannot describe a complete inflationary epoch and should be accompanied with some stable phases. As far as we know, results (2.36) and (2.37) are new.
@@
For solution (2.31), $\beta$ is taken to be exactly zero. In this case, a slight change in initial velocity will result in switching to another constant-roll solution and never returning to $\beta=0$ again. As a result, this solution is also unstable. In general, all $\beta<-3/2$ solutions are unstable, while all $\beta\geq-3/2$ solutions, except for $\beta=0$, are stable.
This work does not focus on studying non-Gaussianity. However, the non-Gaussian properties of solution (2.13) are particularly interesting and worth discussion. In this case, the non-Gaussianity parameter can be derived as follows [11, 12] (see Appendix A):
@@
A look at the evolution of curvature perturbations on supper Hubble scales ($k\ll aH$) can help explain the sudden change of behavior in equations (2.35) and (2.38) at $\beta=-3/2$. In de Sitter limit, the superhorizon curvature perturbation $\mathcal{R}$ is given by (see equation (A.5) for $c_{s}=1$)
$\mathcal{R}=C_{1}+C_{2}\int\frac{\mathrm{~{}d}\tau}{a^{2}\epsilon_{1}}=C_{1}+\frac{C_{2}}{a_{0}^{3}H}\int\frac{\mathrm{~{}d}N}{e^{3N}\epsilon_{1}},$ (2.39)
where $C_{1}$ and $C_{1}$ are two arbitrary constants and $a_{0}=a(N=0)$. As can be seen in equation (2.39), the curvature perturbation has a constant part and an integral part. If $\epsilon_{1}$ is decaying slower than $e^{-3N}$, then the mode function will converge to a constant at super Hubble scales. In contrast, when $\epsilon_{1}$ decays faster than $e^{-3N}$, the integral part diverges and the mode function continues to evolve at large scales. Consistency relation requires that the long modes are well frozen out at large scales, which is not the case here. Therefore, Maldacena’s methodology for proving non-Gaussianity is small does not apply here, and consistency relation cannot be satisfied for fast decaying $\epsilon_{1}$. In equation (2.19), it is evident that for $\beta=-3/2$, the slow-roll parameter decays as $e^{-3N}$. This completely explains why equations (2.35) and (2.38) suddenly change their behavior at $\beta=-3/2$.
@@ 3 Constant-Roll Condition in Non-Canonical Models
In the previous section, the constant-roll condition was defined by three equivalent expressions given in (2.6). In this section, we consider each expression as a distinct constant-roll condition and impose it on a general $P(X,\phi)$ action to show that the nature of the model and its consequences is sensitive to the expression we use. Furthermore, the equivalence is restricted to some models, like single field canonical case, and the result would be different for either expression, in other inflationary models. This comparative study shows how different characteristics of a constant-roll model are contingent upon the constancy of each expression.
We consider a general action as follows:
$S=\frac{1}{2}\int d^{4}x\sqrt{-g}[R+2P(X,\phi)],$ (3.1)
where $P(X,\phi)$ is an arbitrary function of $X$ and $\phi$, and $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. In FLRW background with (,+,+,+)(-,+,+,+)( - , + , + , + ) signature, $X=\dot{\phi}^{2}/2$ corresponds to the canonical kinetic term. The Friedman equations for this action are
@@
Here, $c_{s}$ represents the sound speed which can be expressed as
@@
and the non-trivial field dynamics is controlled by the friction term, $3c_{s}^{2}H\dot{\phi}$, and an effective potential given by
@@
where subscript F refers to a constant-roll condition that constrains the field rate of roll. We call this constraint the field-constant-roll (FCR) condition. This constant-roll condition can be written as $\mathrm{~{}d}\ln\dot{\phi}/\mathrm{~{}d}N=\beta_{\text{\tiny{F}}}$, indicating that the scalar field has a constant roll rate. Integrating it out, we find
@@
As the second approach, we define a constant-roll condition that (similar to the canonical case) leads to a constant $\epsilon_{2}$. Provided that the first slow-roll parameter is small, [^2] This is usual in the standard models of inflation based on inflationary attractor dynamics. then the second slow-roll parameter $\epsilon_{2}=2\epsilon_{1}+\ddot{H}/(H\dot{H})$ is roughly constant by imposing the following condition:
@@
We refer to this constraint as the Hubble-constant-roll (HCR) condition. Note that, unlike the canonical case, $\beta_{\text{\tiny{H}}}$ does not always equal to $\ddot{\phi}/(H\dot{\phi})$. The HCR condition does not ensure a constant rate of roll for the scalar field. Likewise, the FCR condition does not guarantee a constant $\epsilon_{2}$. An interesting property of the HCR condition is that we can directly integrate it to obtain the Hubble parameters given by (2.13)-(2.15), and (2.29). Accordingly, the slow-roll parameters and the acceptable ranges for $N$ and $\beta_{\text{\tiny{H}}}$ are also similar to the canonical case (unless the field becomes imaginary and imposes additional constraints).
@@
where we have defined the effective kinetic term, $Q(X,\phi)\equiv P(X,\phi)+V(\phi)$. From this, we notice that when the FCR condition is imposed, $\ddot{H}/(H\dot{H})$ is not necessarily constant. In a particular model, if the right-hand side of equation (3.10) becomes constant, the HCR condition is met. In this case, we can use (3.9) to rewrite (3.10) as
@@
A key property of the canonical constant-roll model is that it leads to USR for $\beta=-3$, and it can be considered a generalization of the USR model. However, our FCR and HCR conditions do not always correspond to the generalization of USR. In other words, in a particular model, these constant-roll conditions may not reduce to a flat USR potential for any choice of the constant-roll parameter. As our final approach to the generalized constant-roll condition, we define the constant-roll conditions so that the profile of the scalar potential $V(\phi)$ is sufficiently flat when constant-roll parameter approaches $-3$. We approximate the field equation (3.4) in the USR limit as
@@
Although our third manipulation of constant-roll constraint makes implicit use of the third expression in (2.6), it is important to emphasize that condition (3.14) is not the only possible way to generalize the USR. For instance, the condition $\dot{V}=(3+\beta_{\text{\tiny{V}}})HV$, which again exhibits a flat potential when $\beta_{\text{\tiny{V}}}\rightarrow-3$, can be viewed as another generalization of USR and may provide new insights into the model. Nevertheless, the form (3.14) has two advantages: it is identical to FCR and HCR conditions in the canonical limit, and for the examples we will discuss later, $\beta_{\text{\tiny{V}}}\rightarrow 0$ leads to slow-roll case. Indeed, in this work, we discuss only a small number of an apparently endless variety of possible constant-roll conditions picked out by the demand of having a canonical counterpart. [^3] The authors in [28] have studied the constant-roll condition in $f(R)$ gravity by introducing the condition $\ddot{F}=\beta H_{J}\dot{F}$ and $F\equiv df/dR$ to system of equations in Jordan frame. This condition, when transformed to Einstein frame through the relevant transformation relations, depends on the functional form of $f(R_{J})$. Note that all three constant roll conditions in section 3 are independent of the functional form of $P(X,\phi)$; this will let us to talk about some general dynamical behaviors for generic $P(X,\phi)$ and compare the results. The generalization can be done in many ways and with different motivations [28, 40, 20, 24]. In reference [28], for example, simplicity and aesthetic elegance are stated as the reasons for the proposed condition. We concentrated on the conditions with canonical counterpart. In this work, we limit our analysis to $P(X,\phi)$ action, leaving the study of the modified gravity theories for future investigations.
@@
To find the relation between $\beta_{\text{\tiny{V}}}$ and $\beta_{\text{\tiny{H}}}$, we elaborate on equation (3.3) to rewrite the PCR condition (3.13) as
@@
This implies that given a non-canonical action, two constant-roll models based on PCR and HCR constraints would be homologous if we have
@@
Finally, note that the system of equations (3.11), (3.15), and (3.17) are dependent, in the sense that when every pair of the equations is satisfied, all three kinds of constant-roll conditions lead to similar results. The canonical constant-roll model satisfies these conditions for $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}$. In what comes next, we arrange to clearly distinguish between the above approaches for generalizing the constant-roll conditions.
@@ 4 Constant-Roll and Propagation Inflation
This section considers a special class of inflationary models with constant sound speed to demonstrate the applications of the constant-roll conditions discussed in section 3. With constant $c_{s}$, the integration of equation (3.5) gives us
@@
For the class of constant-roll and propagation models, equations (3.2)-(3.4) turn into
@@
$\ddot{\phi}+3c_{s}^{2}H\dot{\phi}+\mathcalboondox{V}_{,\phi}=0.$ (4.4)
Here, the derivative of effective potential is given by
@@
In this section, we study three classes of constant-roll and propagation models constrained with constant-roll conditions introduced in the previous section. However, before proceeding, it is useful to determine whether these classes would be homologous or not. The effective kinetic term in the constant propagation model is given by
@@
Using this, we can show that constant propagation models satisfy equation (3.17) with $2\beta_{\text{\tiny{H}}}=\left(1+c_{s}^{2}\right)\beta_{\text{\tiny{V}}}$. In other words, in a constant propagation model if $\beta_{\text{\tiny{H}}}$ is constant, $\beta_{\text{\tiny{V}}}$ will also be constant and vice versa. Therefore, HCR and PCR constrained classes are not independent and do not lead to different dynamics. By knowing this, we can easily discuss the model’s USR limit. For the HCR condition the Hubble parameter can be derived identical to the canonical case, and since in the constant propagation model the HCR and PCR condition are homologous, the Hubble parameter and slow-roll parameters in the USR case of the constant propagation models are given by equations (2.13) and (2.19), but with $\beta$ replaced with $-3\left(c_{s}^{2}+1\right)/2$. As we discussed in section 2, if $\epsilon_{1}$ decays faster than $e^{-3N}$ modes keep growing at superhorizon scales. Thus, the model cannot be stable unless the condition $\beta=-3\left(c_{s}^{2}+1\right)/2>-3/2$, or equivalently $c_{s}^{2}<0$, is satisfied, which is impossible. It follows that the USR limit of constant propagation model (4.1) can never be stable.
@@
This means that FCR and HCR constrained classes are generally independent and lead to different solutions unless the offending non-universal term ˙$\dot{\mathcal{F}}/(2H\mathcal{F})$ is a constant. Therefore, for a general $\mathcal{F}$ function, we only need to consider one of the HCR and PCR conditions, and we will study the FCR condition separately.